jim has one more than twice as many times than quarters. he has no other coins. if he has $3.25 in change, how many of each coin does jim have.
Answer by ankor@dixie-net.com(22740)

Jim has one more than twice as many dimes than quarters. he has no other coins.
d = 2q + 1
If he has $3.25 in change,
.10d + .25q = 3.25
Replace d with (2q+1)
.10(2q+1) + .25q = 3.25
.20q + .1 + .25q = 3.25
.45q = 3.25 - .10
.45q = 3.25 - .10
q = 3.15/.45
q = 7 quarters
then
d = 2(7) + 1
d = 15 dimes
:
:
Check
.10(15) + .25(7) =
1.50 + 1.75 = 3.25

: